# Filling of incomplete sinograms from sparse PET detector configurations using a residual U-Net

### v2

@@ 1 Introduction
Positron Emission Tomography (PET) imaging has become a cornerstone in medical diagnostics. Technology-driven developments such as iterative image reconstruction and increased computational power have improved performance and paved the way for constructing long axial field-of-view (FOV) "total" body PET systems. Current commercial models like the United Imaging uExplorerTM total-body PET scanner and the Siemens Biograph Vision QuadraTM nearly total-body PET scanner, offering axial FOVs of up to 194 cm and 106 cm, respectively, are examples of this latest technological leap [1, 2]. The updated PennPET Explorer further exemplifies this trend towards extended FOV [3, 4, 5]. These advancements promise benefits like lower radioactive doses, shorter scan times, and reduced image noise. However, they come at a significant cost, primarily due to the increased number of detectors required [6, 7, 8, 9]. Therefore, the traditional cylindrical configuration of PET scanners, characterised by densely packed photodetectors, presents substantial financial limitations in the transition to extended FOV PET imaging.
@@
A possible, cost-effective solution is to introduce gaps between detector elements, either in the axial or transaxial direction or a combination thereof. Such a system geometry potentially extends detector coverage without requiring additional crystal material, thereby providing an extended FOV scanner with reduced material requirements compared to conventional models [8, 10, 11, 12, 13]. However, the sparse detector configurations invariably affect image quality, leading to overall sensitivity loss, increased background variability, and reduced spatial resolution and contrast recovery [8, 14, 15, 16, 17]. Simulation-based comparisons have confirmed that such configurations can reduce system cost but substantially lower sensitivity, especially in designs with aggressive sparsity levels [18, 19]. Nonetheless, cost-effectiveness analyses suggest that extended FOV PET systems may still be financially justifiable in select clinical applications, particularly when paired with algorithmic compensation strategies [20].   Various compensatory methods have been explored to address challenges related to detector gaps, including compressed sensing and inpainting [21, 22, 23], sinogram interpolation [24], optimisation-based reconstruction algorithms [25], normalisation to account for missing lines of response (LORs) [10, 11, 19], and continuous bed motion [9, 17, 26]. However, these solutions have often been applied to small-scale data losses and struggle to handle the massive count losses that arise from removing large portions of detectors elements from the PET configurations. The normalisation approach, although offering a large-scale solution, has previously only been applied to either axial or transaxial gaps of small sizes and has yet to be successfully implemented on the complex missing data patterns resulting from sparsity patterns that combine axial and transaxial gaps. Furthermore, while continuous bed motion helps average out undersampling artifacts, it does not address the root problem of missing detector elements and remains available only on select scanners, limiting its applicability in sparse detector configurations.
@@
In recent years, deep learning has emerged as a potent tool in medical imaging, particularly for data completion tasks like estimating high-count PET data from low-count measurements [27, 28, 29]. A recent review highlights the expanding role of deep learning across the PET reconstruction pipeline [30]. The detector gap problem has been studied as a learning problem in smaller scales such as a few uniformly spaced axial gaps [31, 32, 33, 34], or a C-shaped ring design [35]. However, these studies did not address the extensive and complex patterns of missing data that arise in more aggressive sparse PET configurations, such as those explored here. Moreover, most deep learning applications in PET reconstruction have focused on the image domain; direct sinogram restoration from sparsely sampled data remains largely unexplored.
@@ 2.2 Sinogram restoration network
The proposed deep network was inspired by a segmentation task [36] and consists of a residual U-Net [37] combining the strengths of residual learning [38] and the well-known U-Net architecture [39]. The network consists of an asymmetric encoder-decoder structure (Figure 3). The encoder layers feature an increasing number of residual blocks and one strided convolution, which halves the spatial dimensions. The layers of the decoder are composed of strided transposed convolutions which double the spatial dimensions, followed by one residual block. Skip connections connect the down-sampling and up-sampling segments, allowing the network to combine positional and contextual information to predict the missing sinogram information and aid gradient back-propagation during training. Each sinogram was normalised to the range $[0:1]$ before input and then denormalised by the same factor after output to control the unboundedness of the sinogram values. This was shown to be beneficial during training, achieving a more steady and fast convergence [40]. Finally, the pixels not affected by the removed detectors were copied from the input sinograms and reinstated in the output.
@@ 2.3 Network training
The loss function was minimised using the Adam optimiser [41] for 200 epochs with a learning rate of $10^{-3}$, which experienced a stair-case exponential decay rate of 0.96. During each training epoch, batches of 64 sinograms were used until the whole training set was exhausted. The training was evaluated using the validation set at the end of each epoch. If the validation loss did not improve for 20 consecutive epochs, an early stopping criterion was triggered, effectively stopping the training.
@@ 2.4 Quantification
The sinogram restoration network performance was evaluated in both sinogram and image domains. For comparison, interpolated 2D sinograms were also created using the Clough-Tocher method [42, 43]. Both the network restored (all $1981\times 2$ slices in the test set) and the interpolated data were compared to the ground truth original scans (without missing detectors). Between-sinogram and between-image comparisons were based on the SSIM and MAE in Equations (1)-(2). Note that high values are better for the SSIM, while low values of the MAE are the best.
To complement these global metrics, we also evaluated background variability (BV), defined as the coefficient of variation of ROI mean values, expressed as a percentage:
$BV=\frac{\sigma_{ROI}}{\mu_{ROI}}\cdot 100$
where $\mu_{ROI}$ is the average ROI mean value and $\sigma_{ROI}$ is the standard deviation across ROI mean values. In this study, we report a relative background variability (rBV), defined as the ratio of BV in the restored images to BV in the corresponding original images.
@@ 3.1 Sinogram restoration
The interested reader is referred to supplementary Figure S.1 to inspect the same visual comparison of summed cross-plane sinogram slices with ring difference greater than 1.
@@ 3.2 Reconstructed images
Image slices from the same scans of Figure 7 but reconstructed in coronal view can be found in supplementary Figure S.2.
@@ 4 Discussion
This study presents a novel method for deep learning-based filling of incomplete sinograms resulting from sparse PET configuration systems with inter-detector gaps following a chessboard pattern. Motivated by other studies on sparse detector configurations [8, 9, 10, 11, 14, 15, 16, 17, 24, 26, 25], we explored several patterns for detector cancellation. In addition to the expected sensitivity loss from missing counts relative to non-sparse designs [16, 17], specific image quality degradations have been reported with increasing gap sizes. Background variability and losses in spatial resolution and contrast recovery tend to increase with the axial gap size [8, 9]. Other known drawbacks are noise artefacts due to axial gaps and streaking artefacts due to transaxial gaps [14, 15, 31]. The chessboard pattern used here was chosen to generate maximum detector sparsity (50% of the detectors) with minimal gap size. We investigated the data loss effects of several block sizes for the chess squares. A preliminary evaluation of the sinogram restoration network on configurations of sizes $1\times 1$, $2\times 2$, and $3\times 4$ showed robustness in terms of SSIM and MAE (data not shown). However, only the best-performing (in terms of SSIM and MAE) chessboard pattern size $1\times 1$ was fine-tuned and included in the presented results.
@@
The 2D sinograms corresponding to summed cross planes with ring difference 1 displayed both zeroed-out lines and overall lower counts values in the remaining pixels and, therefore, differ from the ground truth in all sinogram pixels (Figure 2(b)). Hence, the sinogram restoration network had to learn to predict entire sinogram slices for these cross planes. This approach is not possible for the interpolation, where reference points are needed to generate predictions. Therefore, a different mask was supplied to the interpolation method for these mentioned cross planes, leaving the low-count pixels as reference points, resulting in an overall underestimation of the zeroed-out pixels. Even when boosted with a global scaling factor, the interpolation approach consistently underestimated counts and could not match the performance of our restoration network.
The proposed sparse chessboard configuration is designed to explore the potential for lower-cost or more cost-effective clinical PET scanners, aiming for the same axial FOV as conventional scanners but with only 50% of the original detector materials. We further speculate that the proposed sparse geometry could alternatively be used to allow the development of a two-fold extension of the axial FOV of conventional compact ring configurations with the same number of detectors. An extended axial FOV would, for instance, allow simultaneous imaging of distant organs such as the brain and heart, as in the total body PET systems available on the market. This enhanced coverage enables both static and dynamic whole-body imaging, supporting parametric PET – a challenge with standard systems due to axial limitations. A sparse long axial FOV PET scanner with 50% fewer detectors, retaining only 25% of the initial LORs, would inevitably have substantially lower sensitivity than full-coverage total-body scanners. However, its sensitivity would remain comparable to that of earlier-generation PET systems, which have more limited axial coverage. From this perspective, such a system could still enable dynamic whole-body imaging at a reasonable cost and acceptable sensitivity. Combining the two approaches of extended axial FOV and reduced detector cost compared to compact designs results in a medium-cost extended FOV, which offers many of the advantages of a full total body PET but positioning itself between the current conventional PET and extended FOV systems in terms of investments. We note, however, that while the 1×1 sparsity pattern performed best in our evaluation, such a configuration may not be the most practical or cost-effective from a manufacturing perspective. The present work should therefore be regarded as a proof-of-concept from a deep learning standpoint, with practical implementation requiring further study.
@@
Future work should include an investigation of incorporating sharpness to the U-Net structure to mitigate the smoothing effect inherent in neural networks. It would also be interesting to investigate this specific model's capacity regarding possible data losses for clinically approved image quality and if there is an optimal sampling pattern. One might consider an adaptive sparsity pattern, e.g., tighter sampling at the scanner centre and more sparse sampling at the scanner ends. Another necessary step is to explore the integration of sinogram re-normalisation to account for the missing lines of response, particularly in the cross-planes. In this study, we did not explicitly apply re-normalisation. However, our results show that the network is capable of restoring both the missing signals in the direct planes and the attenuated signals in the cross-planes without a separate re-normalisation. This suggests that the model can, to some extent, compensate for the effects of missing and attenuated signals without explicit need for re-normalisation of the sinograms. Nonetheless, applying explicit re-normalisation as a preprocessing step may reduce the network's burden and improve restoration accuracy – a possibility that warrants further investigation.
@@
Furthermore, the dataset was limited in the sense that it only included one anatomical region and only left one patient for validation. This was a main motivation for training the model on individual sinogram slices, which comes at a possible trade-off of introduced bias in the results. The current model was furthermore trained on one specific detector pattern at a time. To increase model generalisability, the dataset should be extended to multiple anatomical regions, and the training process could be extended to include several pattern sizes. Our model could also be combined with parallax correction. In the case of missing LORs due to the missing detector elements, such a correction scheme would also give the added benefit of shifting some of the measured counts to empty detector positions, thereby assisting the estimation of missing counts. Another alternative is to extend our model to include time-of-flight (TOF) information. We expect that including the TOF timing resolution would improve lesion detectability compared to the non-TOF setting due to enhanced signal-to-noise, mainly due to the increase the number of 2D sinograms available for training. We do not expect the reported blurring to go away since the same detector elements are still missing. An additional future direction would be to evaluate whether combining our method with, for instance, bed motion could enhance sinogram restoration, since bed motion effectively provides multiple slightly shifted sampling configurations of the same patient position.
@@
Finally, although our work focuses on sparse ring-based PET configurations, other approaches share the same goal of reducing detector costs through novel system geometries. Flat-panel PET configurations [44, 45] reflect the same broader objective of enabling affordable PET imaging through hardware sparsity. Preliminary studies have applied deep learning-based denoising to flat-panel PET systems with sparse detectors, showing promising image quality with a 30% detector reduction [46].

### v3

@@ 1 Introduction
Positron Emission Tomography (PET) imaging has become a cornerstone in medical diagnostics. Technology-driven developments such as iterative image reconstruction and increased computational power have improved performance and paved the way for constructing long axial field-of-view (FOV) "total" body PET systems. Current commercial models like the United Imaging uExplorerTM total-body PET scanner and the Siemens Biograph Vision QuadraTM nearly total-body PET scanner, offering axial FOVs of up to 194 cm and 106 cm, respectively, are examples of this latest technological leap [uExplorer, Vision]. The updated PennPET Explorer further exemplifies this trend towards extended FOV [PennPET, PennPET2, PennPETex]. These advancements promise benefits like lower radioactive doses, shorter scan times, and reduced image noise. However, they come at a significant cost, primarily due to the increased number of detectors required [Vandenberghe2020, Abgral2021, Daube-Witherspoon2021, Karakatsanis2022]. Therefore, the traditional cylindrical configuration of PET scanners, characterised by densely packed photodetectors, presents substantial financial limitations in the transition to extended FOV PET imaging.
@@
A possible, cost-effective solution is to introduce gaps between detector elements, either in the axial or transaxial direction or a combination thereof. Such a system geometry potentially extends detector coverage without requiring additional crystal material, thereby providing an extended FOV scanner with reduced material requirements compared to conventional models [Daube-Witherspoon2021, Knopp2019, ZeinKarakatsanis2020, Karp2022, Vandenberghe2023]. However, the sparse detector configurations invariably affect image quality, leading to overall sensitivity loss, increased background variability, and reduced spatial resolution and contrast recovery [Daube-Witherspoon2021, Yamaya2008, Yamaya2009, AbiAklVandenberghe2019, ZeinKarakatsanis2021]. Simulation-based comparisons have confirmed that such configurations can reduce system cost but substantially lower sensitivity, especially in designs with aggressive sparsity levels [dadgar2023simulation, gao2023cost_effective]. Nonetheless, cost-effectiveness analyses suggest that extended FOV PET systems may still be financially justifiable in select clinical applications, particularly when paired with algorithmic compensation strategies [alberts2025costeffect].   Various compensatory methods have been explored to address challenges related to detector gaps, including compressed sensing and inpainting [Ahn2012, Valiollahzadeh2015a, Shojaeilangari2018], sinogram interpolation [Karakatsanis2018], optimisation-based reconstruction algorithms [Zhang2019], normalisation to account for missing lines of response (LORs) [Knopp2019, ZeinKarakatsanis2020, gao2023cost_effective], and continuous bed motion [Karakatsanis2022, ZeinKarakatsanis2021, Karakatsanis2019]. However, these solutions have often been applied to small-scale data losses and struggle to handle the massive count losses that arise from removing large portions of detectors elements from the PET configurations. The normalisation approach, although offering a large-scale solution, has previously only been applied to either axial or transaxial gaps of small sizes and has yet to be successfully implemented on the complex missing data patterns resulting from sparsity patterns that combine axial and transaxial gaps. Furthermore, while continuous bed motion helps average out undersampling artifacts, it does not address the root problem of missing detector elements and remains available only on select scanners, limiting its applicability in sparse detector configurations.
@@
In recent years, deep learning has emerged as a potent tool in medical imaging, particularly for data completion tasks like estimating high-count PET data from low-count measurements [Zaidi2020, Zaidi2021, WangDaldrup-Link2023]. A recent review highlights the expanding role of deep learning across the PET reconstruction pipeline [hashimoto2024review]. The detector gap problem has been studied as a learning problem in smaller scales such as a few uniformly spaced axial gaps [Sheikhzadeh2019, Amirrashedi2021, Whiteley2019, LuFeng2021], or a C-shaped ring design [Huang2019]. However, these studies did not address the extensive and complex patterns of missing data that arise in more aggressive sparse PET configurations, such as those explored here. Moreover, most deep learning applications in PET reconstruction have focused on the image domain; direct sinogram restoration from sparsely sampled data remains largely unexplored.
@@ 2.2 Sinogram restoration network
The proposed deep network was inspired by a segmentation task [Myronenko2019] and consists of a residual U-Net [zhang2018road] combining the strengths of residual learning [he2016deep] and the well-known U-Net architecture [U-net]. The network consists of an asymmetric encoder-decoder structure (Figure 3). The encoder layers feature an increasing number of residual blocks and one strided convolution, which halves the spatial dimensions. The layers of the decoder are composed of strided transposed convolutions which double the spatial dimensions, followed by one residual block. Skip connections connect the down-sampling and up-sampling segments, allowing the network to combine positional and contextual information to predict the missing sinogram information and aid gradient back-propagation during training. Each sinogram was normalised to the range $[0:1]$ before input and then denormalised by the same factor after output to control the unboundedness of the sinogram values. This was shown to be beneficial during training, achieving a more steady and fast convergence [lecun2002efficient]. Finally, the pixels not affected by the removed detectors were copied from the input sinograms and reinstated in the output.
@@ 2.3 Network training
The loss function was minimised using the Adam optimiser [reddi2018on] for 200 epochs with a learning rate of $10^{-3}$, which experienced a stair-case exponential decay rate of 0.96. During each training epoch, batches of 64 sinograms were used until the whole training set was exhausted. The training was evaluated using the validation set at the end of each epoch. If the validation loss did not improve for 20 consecutive epochs, an early stopping criterion was triggered, effectively stopping the training.
@@ 2.4 Quantification
The sinogram restoration network performance was evaluated in both sinogram and image domains. For comparison, interpolated 2D sinograms were also created using the Clough-Tocher method [Alfeld1984, Farin1986]. Both the network restored (all $1981\times 2$ slices in the test set) and the interpolated data were compared to the ground truth original scans (without missing detectors). Between-sinogram and between-image comparisons were based on the SSIM and MAE in Equations (1)-(2). Note that high values are better for the SSIM, while low values of the MAE are the best.
To complement these global metrics, we measured mean activity ($\mu_{ROI}$) and standard deviation ($\sigma_{ROI}$) in a 5-10 cm3 bladder region of interest (ROI) and a randomly placed background ROI of similar size. We define background variability (BV) for a ROI as:
$BV=\frac{\sigma_{ROI}}{\mu_{ROI}}\cdot 100\,\%.$
. In this study, we report a relative background variability (rBV), defined as the ratio of BV in the restored images to BV in the corresponding original images. We further evaluated the contrast recovery (CR) defined as the ratio between the contrast in the restored and the original images, defined as:
$CR=\frac{\mu_{\text{bladder,restored}}-\mu_{\text{background,restored}}}{\mu_{\text{bladder,original}}-\mu_{\text{background,original}}}\cdot 100\,\%.$
@@ 3.1 Sinogram restoration
Figure 4 shows representative 2D direct sinograms and summed cross-planes with ring difference 1. The interested reader is referred to supplementary Figure S.1 for corresponding examples from the same scans for larger ring differences (5 and 11), confirming similar cancellation patterns and restoration behaviour across oblique planes.
@@ 3.2 Reconstructed images
Image slices from the same scans of Figure 7 but reconstructed in coronal view can be found in supplementary Figure S.2, supporting the same qualitative trends while revealing some band-shaped artefacts, most evident in the interpolated images.
@@ 4 Discussion
This study presents a novel method for deep learning-based filling of incomplete sinograms resulting from sparse PET configuration systems with inter-detector gaps following a chessboard pattern. Motivated by other studies on sparse detector configurations [Daube-Witherspoon2021, Karakatsanis2022, Knopp2019, ZeinKarakatsanis2020, Yamaya2008, Yamaya2009, AbiAklVandenberghe2019, ZeinKarakatsanis2021, Karakatsanis2018, Karakatsanis2019, Zhang2019], we explored several patterns for detector cancellation. In addition to the expected sensitivity loss from missing counts relative to non-sparse designs [AbiAklVandenberghe2019, ZeinKarakatsanis2021], specific image quality degradations have been reported with increasing gap sizes. Background variability and losses in spatial resolution and contrast recovery tend to increase with the axial gap size [Daube-Witherspoon2021, Karakatsanis2022]. Other known drawbacks are noise artefacts due to axial gaps and streaking artefacts due to transaxial gaps [Yamaya2008, Yamaya2009, Sheikhzadeh2019]. The chessboard pattern used here was chosen to generate maximum detector sparsity (50% of the detectors) with minimal gap size. We investigated the data loss effects of several block sizes for the chess squares. A preliminary evaluation of the sinogram restoration network on configurations of sizes $1\times 1$, $2\times 2$, and $3\times 4$ showed robustness in terms of SSIM and MAE (data not shown). However, only the best-performing (in terms of SSIM and MAE) chessboard pattern size $1\times 1$ was fine-tuned and included in the presented results.
@@
The 2D sinograms corresponding to summed cross planes with ring difference 1 displayed both zeroed-out lines and overall lower counts values in the remaining pixels and, therefore, differ from the ground truth in all sinogram pixels (Figure 2(b)). Hence, the sinogram restoration network had to learn to predict entire sinogram slices for these cross planes. This approach is not possible for the interpolation, where reference points are needed to generate predictions. Therefore, a different mask was supplied to the interpolation method for these mentioned cross planes, leaving the low-count pixels as reference points, resulting in an overall underestimation of the zeroed-out pixels. Even when boosted with a global scaling factor, the interpolation approach consistently underestimated counts and could not match the performance of our restoration network. The effect is likely responsible for the band-shaped artifacts observed in the reconstructed coronal images (Figure S.2). This behaviour may be related to the 2D formulation adopted in this work, where the network was trained on individual sinogram slices without explicit modelling of inter-slice correlations. Further investigation into architectures that better capture cross-plane dependencies may therefore be warranted.
@@
Building on these observations, future work should explore extending the current 2D network into a 3D formulation to explicitly model volumetric context, as well as an investigation of incorporating sharpness to the U-Net structure to mitigate the smoothing effect inherent in neural networks. It would also be interesting to investigate this specific model's capacity regarding possible data losses for clinically approved image quality and if there is an optimal sampling pattern. One might consider an adaptive sparsity pattern, e.g., tighter sampling at the scanner centre and more sparse sampling at the scanner ends. Another necessary step is to explore the integration of sinogram re-normalisation to account for the missing lines of response, particularly in the cross-planes. In this study, we did not explicitly apply re-normalisation. However, our results show that the network is capable of restoring both the missing signals in the direct planes and the attenuated signals in the cross-planes without a separate re-normalisation. This suggests that the model can, to some extent, compensate for the effects of missing and attenuated signals without explicit need for re-normalisation of the sinograms. Nonetheless, applying explicit re-normalisation as a preprocessing step may reduce the network's burden and improve restoration accuracy – a possibility that warrants further investigation.
@@
Furthermore, the dataset was limited in the sense that it only included one anatomical region and only left one patient for validation. This was a main motivation for training the model on individual sinogram slices, which comes at a possible trade-off of introduced bias in the results. The current model was furthermore trained on one specific detector pattern at a time. To increase model generalisability, the dataset should be extended to multiple anatomical regions, and the training process could be extended to include several pattern sizes. Our model could also be combined with parallax correction. In the case of missing LORs due to the missing detector elements, such a correction scheme would also give the added benefit of shifting some of the measured counts to empty detector positions, thereby assisting the estimation of missing counts. Another alternative is to extend our model to include time-of-flight (TOF) information. We expect that including the TOF timing resolution would improve lesion detectability compared to the non-TOF setting due to enhanced signal-to-noise, mainly due to the increase in the number of 2D sinograms available for training. We do not expect the reported blurring to go away since the same detector elements are still missing. An additional future direction would be to evaluate whether combining our method with, for instance, bed motion could enhance sinogram restoration, since bed motion effectively provides multiple slightly shifted sampling configurations of the same patient position. In summary, the present work should be regarded as a proof-of-concept from a deep learning standpoint, with practical implementation requiring further development.
The proposed sparse chessboard configuration is designed to explore the potential for lower-cost or more cost-effective clinical PET scanners, aiming for the same axial FOV as conventional scanners but with only 50% of the original detector materials. We further speculate that the proposed sparse geometry could alternatively be used to allow the development of a two-fold extension of the axial FOV of conventional compact ring configurations with the same number of detectors. An extended axial FOV would, for instance, allow simultaneous imaging of distant organs such as the brain and heart, as in the total body PET systems available on the market. This enhanced coverage enables static whole-body imaging, a challenge with standard systems due to axial limitations. A sparse long axial FOV PET scanner with 50% fewer detectors, retaining only 25% of the initial LORs, would inevitably have substantially lower sensitivity than full-coverage total-body scanners. However, its sensitivity would remain comparable to that of earlier-generation PET systems, which have more limited axial coverage. From this perspective, such a system could still enable whole-body imaging at a reasonable cost and acceptable sensitivity. Combining the two approaches of extended axial FOV and reduced detector cost compared to compact designs results in a medium-cost extended FOV, which offers many of the advantages of a full total body PET but positioning itself between the current conventional PET and extended FOV systems in terms of investments. We note, however, that while the 1×1 sparsity pattern performed best in our evaluation, such a configuration may not be the most practical or cost-effective from a manufacturing perspective.
@@
Finally, although our work focuses on sparse ring-based PET configurations, other approaches share the same goal of reducing detector costs through novel system geometries. Flat-panel PET configurations [vandenberghe2023flatpanel, abi2025flatpanel] reflect the same broader objective of enabling affordable PET imaging through hardware sparsity. Preliminary studies have applied deep learning-based denoising to flat-panel PET systems with sparse detectors, showing promising image quality with a 30% detector reduction [abiakl2024deep-learning].
@@ Acknowledgments
Supplementary Materials
Figure S.1: Visual comparison of 2D sinograms based on the ground truth (Original), the sparse chessboard configuration (Distorted), and predictions from interpolation filling (Interpolated) and from the sinogram restoration network (Restored). The figure includes two different sinogram slices from one scan in the test set. (a) shows a sinogram slice with ring difference 5, and (b) shows a sinogram slice with ring difference 11.
Figure S.2: Visual comparison of reconstructed image slices based on original sinograms (Original), sinograms from the sparse chessboard configuration (Distorted), and predicted sinograms from interpolation (Interpolated) and from the restoration network (Restored). (a) and (b) show two different image slices from two different scans in the test set, with reconstructed images in the coronal view on the top vertical panel, corresponding difference images on the middle panel (colour bar in Bq/ml), and corresponding histograms of pixel differences in the bottom panel ($x$-axis in Bq/ml).